### v1

# The Impact of Domain Knowledge and Multi-Modality on Intelligent Molecular Property Prediction: A Systematic Survey

Abstract
The precise prediction of molecular properties is essential for advancements in drug development, particularly in virtual screening and compound optimization. The recent introduction of numerous deep learning-based methods has shown remarkable potential in enhancing molecular property prediction (MPP), especially improving accuracy and insights into molecular structures. Yet, two critical questions arise: does the integration of domain knowledge augment the accuracy of molecular property prediction and does employing multi-modal data fusion yield more precise results than unique data source methods? To explore these matters, we comprehensively review and quantitatively analyze recent deep learning methods based on various benchmarks. We discover that integrating molecular information will improve both MPP regression and classification tasks by upto 3.98% and 1.72%, respectively. We also discover that the utilizing 3-dimensional information with 1-dimensional and 2-dimensional information simultaneously can substantially enhance MPP upto 4.2%. The two consolidated insights offer crucial guidance for future advancements in drug discovery.
1 Introduction
The field of drug development has always been at the forefront of adopting innovative scientific techniques to enhance the discovery and optimization of therapeutic compounds. Central to this process is the prediction of molecular properties, a task that bears significant implications for drug screening and compound optimization[1]. Accurately predicting key molecular properties can significantly reduce the time and resources required in drug development, thereby hastening the journey towards innovative medical treatments.
In the landscape of computational methods for molecular property prediction (MPP), deep learning (DL) has recently emerged as a transformative force, distinguishing itself markedly from traditional techniques such as quantitative structure-activity relationships (QSAR) and molecular dynamics simulations. While conventional methods have laid the groundwork, DL significantly advances accuracy and analysis depth, enabling a more intricate exploration of the relationships between molecular structures and their properties[2].
Despite the advancements in DL for MPP, the field continues to face ongoing evolution and challenges. Two significant trends are currently shaping the field. The first trend is the increasing integration of domain knowledge into DL models. This includes a broad spectrum of scientific information, such as chemical and physical property relation, atom and bond characteristics, and detailed insights into functional groups and molecular fragments.
Figure 1: The overview of our survey. we review the impact of domain knowledge and multi-modality on molecular property prediction from three critical aspects: input data, model architectures, and training strategy. The detailed information are explained in the following sections.
The integration of this knowledge aims to enhance the predictive accuracy of these models. This leads us to the critical question: Does more comprehensive domain knowledge actually improve the effectiveness of MPP? The second trend is the rising adoption of multi-modality techniques, which involve the fusion of various data types like sequence-based, graph-based and pixel-based formats. This approach is driven by the goal of achieving more accurate predictions in a field characterized by its complexity and data diversity, prompting the question: Is multi-modality more effective for MPP than methodologies that rely on uni-model data source? To explore these questions, our paper begins with an in-depth review of the current DL approaches in MPP, focusing on how the domain knowledge and multi-modal data integrate on encoder architecture and training strategy.
Our review begins with an examination of various unique data encoder architectures for MPP such as Recurrent Neural Network (RNN)[3, 4, 5], Graph Neural Network (GNN)[6, 7, 8, 9, 10], Transformer[11, 12, 13, 14], and Convolutional Neural Network (CNN)[15, 16, 17] models, and also review the multi-modal methods[18, 19, 20, 21] We focus on how these architectures are aligned with existing molecular structural knowledge and their integration of domain knowledge. This exploration highlights the synergy between advanced computational techniques and fundamental molecular understanding, a crucial aspect in enhancing the accuracy of MPP. Also we review a variety of training strategies, such as self-supervised[22, 23, 24], semi-supervised[25, 26, 27], transfer learning[28, 29, 30] and multi-tasks learning[31, 32]. A particular emphasis is placed on strategies that effectively utilize unlabeled data, a vital consideration given the frequent scarcity of labeled data in this domain, and we focus on how the domain knowledge and multi-modal data to be used in the training strategy. Accompanying this review are comprehensive diagrams that systematically elucidate the nuances of these encoder architectures and training strategies, offering a clearer understanding of their complex mechanisms. The overview of our paper is as Figure 1.
Our study then proceeds to empirically evaluate these DL methods, utilizing pivotal benchmarks like MoleculeNet[33]. These benchmarks, encompassing a diverse range of datasets each focused on specific molecular properties, allow for an extensive assessment of different DL approaches. A key aspect of our analysis is determining the impact of multi-modality techniques versus single modeling. Specifically, we investigate the effectiveness of integrating atom-bond level domain knowledge and substructures, such as functional groups and fragments, into the models. Additionally, we quantify the contributions of different data formats and conduct experiments to ascertain whether multi-modal fusion can enhance the generalization performance of the models. This evaluation not only provides comparative insights into the varied methods but also seeks to pinpoint essential factors that bolster the efficacy of DL in MPP.
In summary, our main contributions are as follows:
$\bullet$ We identify two pivotal issues when applying DL for MPP: domain knowledge integration and multi-modal data utilization.
$\bullet$ We comprehensively review DL methods for MPP, featuring in-depth analyses of encoder architectures and training strategies.
$\bullet$ We discover that integrating molecular substructure information results in a 3.98% improvement on average in regression tasks and a 1.72% increase on average in classification tasks.
$\bullet$ We discover that the utilizing 3-dimensional information with 1-dimensional and 2-dimensional information simultaneously can substantially enhance MPP upto 4.2%.
2 Molecular Data
In the field of molecular science, an vast range of molecular data has been developed, each crucial for computational modeling and analysis. These formats are generally classified into three main types: text-based, graph-based, and pixel-based formats. Each type offers unique insights into molecular structures, contributing significantly to various aspects of molecular analysis. These diverse formats are illustrated in the Figure 2, which showcases the array of molecular data available for DL methods.
2.1 Sequence-based Data
Text-based formats are among the most commonly used representations in MPP due to their simplicity and efficiency. The most prominent of these is Simplified Molecular Input Line Entry System (SMILES)[34], which encodes molecules in linear strings, representing atoms and bonds in a compact, readable format. Variants of SMILES, such as Canonical SMILES[35] and Isomeric SMILES[36], offer additional specificity, including stereochemistry information. Other notable text-based formats include molecular fingerprints like ECFP[37], Morgan, and MACCS[38], which encode the presence of certain molecular features, and Self-Referencing Embedded Strings (SELFIES)[39], a newer format designed for robustness in machine learning applications. Additionally, IUPAC[40] and InChI[41] codes are vital text-based molecular representations. IUPAC provides systematic chemical nomenclature for clear scientific communication, while InChI offers standardized textual identifiers for chemical substances. These formats facilitate various computational tasks, from database searching to the generation of novel molecules using AI.
2.2 Graph-based Data
In drug discovery, graph-based representations, which depict atoms as nodes and bonds as edges, effectively capture molecular structures, making them ideal for analyzing both topological and relational aspects of molecules. The method includes the use of a 2D adjacency matrix or a set of edges to outline atom connectivity. This representation can be enhanced with 3D information, such as bond lengths and atom positions, transforming it into a 3D graph. Incorporating a 3D atom distance matrix further enriches this model, offering a comprehensive view of the molecular spatial structure. Graph-based formats, including 2D and 3D molecular structures, are crucial in drug discovery for conducting detailed molecular analyses and enhancing the understanding of complex molecular behaviors.
2.3 Pixel-based Data
Pixel-Based Molecular Data Formats, such as 2D images and 3D grids, are essential components of molecular property prediction. These formats, easily generated by tools like RDKit[42] and PyMol[43] for 2D images and Libmolgrid[44] for 3D grids, offer clear and comprehensible visual representations of molecular structures. This visual aspect allows for straightforward human interpretation, aiding in the recognition of molecular patterns and the understanding of spatial relationships in computational modeling. This clarity in visualization is crucial for effectively analyzing molecular geometries and interactions.
Figure 2: Molecular data: We illustrate the transformation of molecular data formats crucial for MPP using the example of the caffeine molecule. This is demonstrated across three primary categories: sequence-based, graph-based, and pixel-based formats. Each format is derived from the SMILES representation of caffeine, using Python packages such as RDKit and software tools like PyMol. a). The sequence-based data section includes formats like SMILES, its variants (Canonical and Isomeric SMILES), molecular fingerprints (ECFP, Morgan, MACCS), and SELFIES, highlighting their roles in encoding molecular structures. b). Graph-based data represents caffeine as a graph with atoms as nodes and bonds as edges, enriched with 3D information for detailed structural insights. c). Pixel-based data showcases 2D images and 3D grids of caffeine, crucial for visual analysis and spatial interpretation.
3 Domain Knowledge
In molecular science, many domain knowledge from areas like physics, chemistry, and biology plays a vital role. This knowledge is methodically grouped into four key categories: atom-bond property, molecular substructure, chemical reactions, and molecular characteristics. Each category is integral for a comprehensive understanding and accurate interpretation of molecular data. The Figure 3 showcases these categories in detail, providing an in-depth look at the essential aspects of molecular information interpretation.
3.1 Atom-bond Property
In MPP, a deep understanding of atomic and bonding attributes is vital for accurately modeling molecular behaviors. Understanding atomic properties such as isotope numbers, which affect molecular weight and stability, and chirality, crucial for interaction and reaction in biological systems, is essential. Hybridization types impact bonding patterns and molecular geometry. Atomic valence, number, mass, formal charge, and aromaticity significantly influence a molecule’s chemistry.
Figure 3: Molecular domain knowledge: This figure categorizes molecular expert knowledge essential for MPP into four domains, using the molecule C=(CC(=C)c1cc(C(=O)O)cnc1C(C)$C$ as an example. a). In the atom-bond property section, we examine aspects such as the molecule’s atomic number, mass, valence, and bond types, focusing on isotope numbers that influence molecular weight and stability, and chirality affecting how the molecule interacts in biological systems. b). Molecular substructure includes the functional groups, molecular fragments, and pharmacophores of this molecule, illustrating their influence on its chemical behavior and interactions. c). Molecular property covers a range of properties from quantum mechanics to physiology, showcasing how these properties affect the molecule’s behavior in drug development. d). Chemical reaction discusses the mechanisms of molecular transformations, highlighting the molecule’s reactivity.
Bond attributes like bond type and stereochemistry are key in determining molecular connectivity, reactivity, and shape, influencing interactions with biological targets. The direction and length of bonds also provide insights into spatial arrangement. These detailed atomic and bond attributes collectively provide a comprehensive framework for molecular structure analysis, essential for effective predictive modeling in drug discovery.
3.2 Molecular Substructure
In the realm of MPP, a deep comprehension of molecular substructures is indispensable. These substructures, including functional groups, molecular fragments, and pharmacophores, are fundamental in dictating the functions and interactions of molecule.
These substructures, such as functional groups, molecular fragments, and pharmacophores, play key roles in understanding a molecule’s behavior. Functional groups, such as hydroxyl (-OH) and carboxyl (-COOH), are specific group of atoms within a molecule that is responsible for the characteristic chemical reactions of that molecule, and are particularly influential in determining a molecule’s chemical behavior and interactions. For example, a hydroxyl group can significantly increase water solubility, thereby impacting a drug’s absorption, distribution, and overall pharmacokinetics.
Molecular fragments are larger portions of molecules, encompassing various structural elements like rings or chains. Similarly, molecular fragments like benzene rings affect a molecule’s stability and electronic, which in turn can alter its interaction with biological receptors or enzymes, impacting biological activity. Common molecular fragment methods are breaking of retrosynthetically interesting chemical substructure (BRICS)[45], Retrosynthetic Combinatorial Analysis Procedure (RECAP)[46], Murcko scaffolds[47], eMolFrags[48], and rdScaffoldNetwork[49].
A pharmacophore is an abstract representation of the molecular features necessary for a molecule to interact with a specific biological target to produce a desired biological effect. A pharmacophore with both a hydrogen bond donor and an acceptor in a specific spatial arrangement, for instance, can be crucial for binding to biological targets like enzymes or receptors, influencing the molecule’s effectiveness as a therapeutic agent. The accurate identification and understanding of these substructures are key to developing new pharmaceuticals, offering detailed insights into molecular interactions.
3.3 Chemical Reaction
Chemical reactions involve the transformation of substances through the breaking and forming of chemical bonds, leading to the creation of new molecules with specific properties. For example, in the reaction C=CC(=C)c1cc(C(=O)O)cc(NC(=O)C)c1C(C)CC + SOCl2 → C=CC(=C)c1cc(C(=O)O)cnc1C(C)CC + HCl + SO2, the reactant interacts with thionyl chloride, resulting in a new product plus byproducts. This process highlights the role of reactants, products, and catalysts in affecting reaction outcomes and mechanisms. Such knowledge is vital for predicting reaction paths, designing new molecules with desired properties, and developing effective pharmaceuticals and novel compounds.
3.4 Molecular Property
MPP in drug discovery is a multidisciplinary field, each discipline offering detailed insights into molecular behavior. Quantum mechanics, for example, delves into electronic properties like ionization potentials, crucial for understanding reaction mechanisms. Physical chemistry examines molecules’ stability, reactivity, and phase behaviors, impacting drug formulation. Biophysics explores molecular interactions within biological systems, crucial for drug-target binding studies. Physiology, on the other hand, assesses drug effects at an organismal level, influencing pharmacodynamics and pharmacokinetics. These interconnected properties, such as how a drug’s solubility impacts absorption and bioavailability, highlight the need for a comprehensive understanding across levels, from atomic to organismal, to predict molecular properties accurately and develop effective pharmaceuticals. This integrative approach, encompassing everything from electron distribution to organismal response, is vital in the nuanced field of drug development.
4 Modeling Method
Our paper provides a concise yet comprehensive examination of current DL methods in MPP. We first review molecular encoder architectures, and explore how these encoder align with the prior structural knowledge of molecules and how domain knowledge is integrated into them. Our review further emphasizes the utilization of unlabeled data, encompassing an exploration of self-supervised, semi-supervised, transfer learning, and multi-task learning strategies. To aid in understanding these intricate concepts, our paper includes detailed diagrams, which elucidate these advanced computational methods and their integration with fundamental molecular insights, thereby contributing to the advancement of MPP.
4.1 Encoder
In MPP, encoder architectures play a key role in transforming raw molecular data into meaningful representations. This section examines a variety of encoder architectures, each tailored to specific molecular data formats and complexities. We categorize four main types of encoders for single data sources: RNN-based, GNN-based, Transformer-based, and CNN-based. Each type is analyzed for its alignment with molecular prior structural knowledge and the integration of domain-specific information. Additionally, we examine multi-modality based encoders, which handle multiple data sources, highlighting their unique characteristics, applications, and the challenges they address in molecular representation learning. The detailed aspects of these encoder architectures are illustrated in the Figure 4.
4.1.1 RNN-based
Figure 4: Molecular encoder architectures: This figure categorizes molecular encoder architectures into four types: RNN-based, Transformer-based, GNN-based, and CNN-based. Each type is assessed for its ability to capture information about functional groups like carboxyl groups (-COOH) in the molecule C=(CC(=C)c1cc(C(=O)O)cnc1C(C)$C$. a). RNN-based encoders process sequence data, maintaining a memory of previous inputs to effectively capture sequential patterns of (-COOH). b). Transformer-based models utilize Self-Attention mechanisms, enabling them to identify and focus on the (-COOH) group’s specific interactions within the molecular sequence. c). GNN-based architectures employ a message passing strategy, extracting the topological information of (-COOH) within the molecule’s graph structure. d). CNN-based models analyze spatial patterns through convolution layers, identifying sub-images that contain the (-COOH) group. This visualization highlights how each encoder type uniquely processes and interprets the molecular structure for MPP.
RNN, like long short-term memory (LSTM)[50] and gated recurrent unit (GRU)[51], are adept at processing sequential data, with a unique internal memory feature that allows them to maintain context and order in sequences. This capability makes RNNs highly effective for tasks involving sequences data. Nowaday some work use RNN-based model to analyze 1D molecular data such as SMILES.
Lin et al.[4] first transformed SMILES into sample vectors, which were then processed using bidirectional GRU neural networks to predict molecular properties, illustrating an innovative approach in training models for molecular property prediction. Lv et al.[5] introduced Mol2Context-vec to address the challenge of representing molecular substructures and their polysemous nature, integrating different internal state levels for dynamic representations. To highlight the SMILES characters that are more important for the prediction tasks, Wu et al.[52] utilized the bidirectional long short term memory attention network in which they employed a novel multi-step attention mechanism to facilitate the extracting of key features from the SMILES strings. Nazarova et al.[53] used the single-layer Elman RNN to identify correlations between the structure of polymers of the norbornene class and their permittivity while using the SMILES notation in binary and decimal representations. Wang et al.[54] employs a Tree-structured LSTM network with signature descriptors to automatically generate expressive signatures for molecular structures, enabling the efficient representation of their structural information and connectivity in a single-step process.
These works demonstrate the effectiveness of RNN in extracting semantic information from SMILES sequences, paralleling methods in natural language processing(NLP). However, they face challenges when incorporating varied expert knowledge and managing long SMILES sequences, and the focus of RNN-based models on adjacent characters hampers effective interactions between distant atoms. This limitation can affect their ability to capture extensive structural relationships, especially when important atoms within the same functional group are distantly placed in the sequence.
4.1.2 GNN-based
Molecules can be effectively represented as graphs, with atoms as nodes and chemical bonds as edges. GNNs are well-suited to learn from this representation, utilizing layers that enable message passing. In GNNs, node embeddings are updated by aggregating information from neighboring nodes, allowing the network to capture molecular features through atom-level interactions. This method provides a detailed understanding of molecular structures by considering both individual atomic characteristics and their interconnections within the molecule. Yang et al.[55] construct molecular encodings by using convolutions centered on bonds instead of atoms, thereby avoiding unnecessary loops during the message passing phase of the algorithm. AttentiveFP[10] not only characterizes the atomic local environment by propagating node information from nearby nodes to more distant ones but also allows for nonlocal effects at the intramolecular level by applying a graph attention mechanism. Withnall et al.[56] introduce attention and edge memory schemes to the existing message passing neural network framework. To address insufficient bond information extraction, Li et al.[57] explicitly drop the matrix mapping of edge features and employ a triplet message mechanism. This mechanism calculates messages from atom-bond-atom information and updates the hidden states of neural networks. Zhang et al.[58] propose CoAtGIN, which uses k-hop convolution to capture long-range neighbor information at the local level and utilizes linear attention to aggregate the global graph representation according to the importance of each node and edge at the global level.
But these methods focus on atom (node) or bond (edge) information. To address this issue, Song et al.[8] propose a Communicative Message Passing Neural Network to improve molecular embedding by strengthening the message interactions between nodes and edges through a communicative kernel. SC-NMP[59] aggregates the node representations of the current step and the graph representation of the previous step, and proposes densely self-connected neural message passing, which connects each layer to every other layer in a feed-forward fashion. To extract useful interactions between a target atom and its neighboring atomic groups, Li et al.[60] propose a new graph learning paradigm based on a block design named block-based GNN and demonstrate that the network degradation problem can be reduced by applying block design with normalization and skip-connection. Ma et al.[61] employ cross-dependent message passing strategy to integrate the node-centered and edge-centered encoders. Liu et al.[62] develop a hypergraph-based topological framework to characterize detailed molecular structures and interactions at the atomic level. They have recently proposed embedding homology and persistent homology. Feng et al.[63] transform each molecular graph into a heterogeneous atom-bond graph to fully utilize the bond attributes and design unidirectional position encoding for such graphs. Biswas[31] pass additional atomic and molecular features, including 2D RDKit descriptors, Abraham parameters, QM descriptors, and 3D geometries, to improve the model performance.
{forest}
for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center,, text width=2.7cm, s sep=6pt, calign=edge midpoint, text=black, , [ Encoder
Architecture , fill=gray!45, text=black, parent [ RNN-based, text=black, rnn [ LSTM-based, rnn_more [ Mol2Context-vec [5]; Wu et al. [52], rnn_work ] ] [ GRU-based, rnn_more [ Lin et al. [4];, rnn_work ] ] ] [ GNN-based, text=black, gnn [ Focusing Topological Structure, gnn_more [ D-MPNN [55]; Attentive FP [10], gnn_work ] ] [ Substructure Enhance, gnn_more [ HimGNN [6]; MPMol [64], gnn_work ] ] [ 3D Information Enhance, gnn_more [ GEM [65]; Uni-Mol [66], gnn_work ] ] ] [ Transformer-based, text=black, transformer [ For Sequence Data, transformer_more [ MolFormer [12]; ChemBERTa [67]; SELFormer [68], transformer_work ] ] [ For Graph Data, transformer_more [ GROVER [11]; LGI-GT [13], transformer_work ] ] ] [ CNN-based, text=black, cnn [ For 2D Image, cnn_more [ ABC-Net [69], cnn_work ] ] [ For 3D Grid, cnn_more [ MR-3D-DenseNet [17];, cnn_work ] ] ] [ Multi-Modality-based, text=black, multi [ Transformer-M [14]; MoleculeSTM [70], multi_work ] ] ]
Figure 5: The molecular encoder architecture summary.
Hasebe[71] proposed a knowledge-embedded message passing nerual network that can be supervised together with nonquantitative knowledge annotations by human experts on a chemical graph. This graph contains information on the important substructure of a molecule and its effect on the target property. Yang et al.[72] extract physical information with a neural physical engine that learns molecular conformations by simulating molecular dynamics with parameterized forces. They then employ this physical information as supplementary data for predicting molecular properties.
However, most methods essentially attribute predictions to individual nodes, edges, or node features. This kind of interpretability is only partially compatible with chemists’ intuition at best. Chemists are more accustomed to comprehending the causal relationship between molecular structures and properties in terms of chemically meaningful substructures, such as functional groups, rather than individual atoms or bonds. Zang et al.[73] decompose the molecular graph by BRICS and additional decomposition to construct a motif-level graph, in which corresponding multi-level generative and predictive tasks are designed as self-supervised signals. As the graph pooling technique for learning expressive graph-level representation is critical yet still challenging, Liu et al.[74] propose master-orthogonal attention, a novel cross-level attention mechanism specifically designed for hierarchical graph pooling. To fully explore higher-order substructure information, Gao et al.[75] propose substructure interaction attention, which takes both the information of neighbors’ substructures and the interaction information among them into account during the aggregation process. To retain locality and linear network complexity, Bouritsas et al.[7] employ a topologically-aware message passing scheme based on substructure encoding, which does not attempt to adhere to the Weisfeiler-Leman hierarchy. Addressing the oversmoothing problem in multi-hop operations, Ye et al.[76] construct a composite molecular representation with multi-substructural feature extraction and process such features effectively with a nested convolution plus readout scheme to capture interacting substructural information. Zhu[77] utilize corepresentation learning of molecular graphs and chemically synthesizable BRICS fragments. Furthermore, a plug-and-play feature-wise attention block is first designed in the their model architecture to adaptively recalibrate atomic features after the message passing phase. To accurately model the complex quantum interactions inherent in molecules, Lu et al.[78] utilize a sophisticated hierarchical graph neural network, which directly extracts features from both the conformation and spatial information of molecules, and then integrates these features through multilevel interactions. Fey et al.[79] take in two complementary graph representations: the raw molecular graph representation and its associated junction tree, where nodes represent meaningful clusters in the original graph. Focusing on the molecular hierarchical relationship, Han et al.[6] propose a simple yet effective rescaling module, called contextual self-rescaling, that adaptively recalibrates molecular representations by explicitly modeling interdependencies between atom and motif features. Ji et al.[64] model a molecule as a heterogeneous graph and leverage metapaths to capture latent features for chemical functional groups. They also design a hierarchical attention strategy to aggregate heterogeneous information at both the node and relation levels. To extract functional groups as motifs for small molecules, Wu et al.[80] construct a heterogeneous molecular graph with both atom-level and motif-level nodes and adopt a heterogeneous self-attention layer to distinguish the interactions between multi-level nodes.
Since different 3D structures may lead to dissimilar molecular properties despite having the same 2D molecular topology, Recently, many works utilizing molecular 3D structures have been introduced. To emphasize equivariant constraints, Fuchs et al.[81] utilize the explicit increase of equivariance constraints in self-attention mechanisms. As rotation-invariant representations struggle to convey directional information, Schutt et al.[82] proposed rotationally equivariant message passing, exemplified by the Polarizable Atom Interaction Neural Network architecture. Brandstetter et al.[83] expand equivariant graph networks to include not only invariant scalar attributes but also covariant information like vectors or tensors. This model consists of steerable MLPs, capable of incorporating geometric and physical information within its message and update functions. Gasteiger et al.[84] show the universality of spherical representations and employ a two-hop message passing mechanism with directed edge embeddings for rotationally equivariant predictions, and utilize symmetric message passing, augmented with geometric information, to enhance our model’s efficacy in MPP. Gasteiger et al.[85] integrate directional information and interatomic distances by embedding and updating messages between atoms, using a spherical 2D Fourier-Bessel basis to jointly represent distances and angles. To model angular relationships among neighboring atoms in a GNN, ensuring constraints like rotation invariance and energy conservation, Shuaibi et al.[86] utilize a per-edge local coordinate frame and innovate a spin convolution, thereby securing rotation invariance in edge messaging. Fang et al.[65] proposed a self-supervised framework using molecular geometric information by constructing a new bond angle graph, where the chemical bonds within a molecule are considered as nodes and the angle formed between two bonds is considered as the edge.
The GNN-based model section concludes by recognizing that while GNN excel in capturing molecular topological information and integrating domain knowledge, their effectiveness is hindered by the small-world phenomenon. This characteristic leads to over-smoothing in deeper networks, where nodes lose feature distinctiveness, impacting predictive accuracy. Additionally, the specialized structure of GNN makes it challenging to scale up with increased parameters, limiting their capability to handle large molecular datasets effectively.
4.1.3 Transformer-based
Originally excelling in NLP, the Transformer architecture is renowned for its self-attention mechanism, which allows for parallel processing of entire sequences. This capability enables it to efficiently manage long-range dependencies within data, making it highly effective in MPP. Its adeptness at understanding detailed contextual relationships enhances the accuracy and computational efficiency in predictive modeling. Wang et al.[87] and Chithrananda et al.[67] use Transformer to extract molecular information from SMILES, which is treated as natural language. Wang et al.[88] proposed two significant advances in molecular data processing: structural fingerprint tokenization for more efficient molecule graph tokenization and normalized graph raw shortcut-connection to enhance latent representations in complex model structures. To address challenges in the validity and robustness of SMILES representations, Yüksel et al.[68] uniquely utilizes SELFIES, a robust and flexible molecular representation format, to learn high-quality molecular features, enhancing the reliability of molecular data analysis in computational chemistry. To predicting activity coefficients in binary mixtures, Winter et al.[89] integrate information from two SMILES strings representing the mixture components, along with temperature and token position data, into a unified matrix for input encoding. Ross et al.[12] explore the diference between absolute and relative position embeddings in representing molecular SMILES and provide a linear attention approximation of the recently proposed relative position RoFormer[90]. Ross et al.[12] delves into the differences between absolute and relative position embeddings in SMILES representation, proposing an efficient linear attention approximation for the RoFormer[90] model, which focuses on relative positioning, to enhance molecular SMILES processing in deep learning applications.
The Transformer architecture, originally designed for sequence data, has been effectively adapted for molecular graph representation in recent research. Its proficiency in handling global molecular information enhances its utility in molecular property prediction, showcasing its versatility beyond traditional sequence analysis. Maziarka et al.[91] proposed the molecule attention transformer, which adapts the Transformer architecture, augmenting the self-attention mechanism with inter-atomic distances and molecular graph structure. Li et al.[9] focus on chemical bonds in molecular representations, employing molecular line graphs to illustrate edge adjacencies in original molecular graphs. Each graph is augmented with a knowledge node containing molecular descriptors and fingerprints, connected to its original nodes Rong et al.[11] combines message passing networks with a transformer-style architecture, extract vectors as queries, keys and values from nodes of the graph, then feed them into the attention block. Yin et al.[13] developed a method that alternates between GNN and Transformer layers, repeated in sequence. This approach effectively blends local and global information, allowing the graph transformer to comprehensively integrate node data from both nearby and distant sources. To extrace the coarse-grained view, Ren et al.[92] make the molecular graph first enters the message passing phase of the traditional GNN layers to update the node embeddings, then enters graph transformation layers to learn different granular information. To achieve one encoder for extracting 2D or 3D information, Luo et al.[14] use two separated channels to encode 2D and 3D structural information and incorporate them with the atom features in the network modules. To fully leverages chemical knowledge, Gao et al.[93] construct an embedding unit comprising a GNN and a transformer to balance the neighboring and distant interactions of an atom, and more attention is given to conjugated systems, unsaturated bonds, heteroatoms and the molecular topology. To extract the molecular fragment information, Jiang et al.[94] design a pharmacophoric-constrained multi-views molecular representation graph, enabling PharmHGT to extract vital chemical information from functional substructures and chemical reactions.
Transformer have demonstrated effectiveness in recent work, particularly with sequence data like SMILES, where they treat it similarly to natural language. Their global information extraction capabilities also extend to molecular graph representation. Recent innovations combine transformer with GNN, enabling simultaneous local and global data analysis. This blend showcases transformer’s strength in handling large molecular datasets and extracting comprehensive insights, vital in MPP.
4.1.4 CNN-based
CNN, known for processing grid-like topology data, are adept at extracting features through convolutional layers and efficiently detecting local patterns. This makes them highly effective for image and pattern recognition tasks, a trait utilized extensively in MPP.
To extract the local pattern of 1D molecular data, Hirohara et al.[95]’s innovative application of CNNs to SMILES data for chemical motif detection marked a significant step in computational drug discovery. Chen et al.[16] highlighted the impact of SMILES molecular enumeration on CNNs’ performance in solubility prediction.
As DL methods have achieved great success in the image processing field, some work used CNN to extracted 2D molecular image, but the size of the same atom/structure is vibrational in different molecules because of the fixed size of the whole molecular image. To address this issue, Zhang et al.[69] introduced ABC-Net, predicting graph structures by representing atoms and bonds as points, utilizing CNN-generated heat-maps. Jiang et al.[96] proposed an equal-sized molecular persistent spectral image, and encoder it with CNN model to extract molecular representation.
As the visual representation of molecular structure, 3D molecular grid is important for extracting molecular 3D information. However, a direct 3D representation of a molecule with atoms localized at voxels is too sparse, which leads to poor performance of the CNNs. To addrress this issue, Denis et al.[97] present a novel approach where atoms are extended to fill other nearby voxels with a transformation based on the wave transform. Shuai et al.[17] utilize an atom-centered gaussian density model for 3D molecular representation, which involves defining multiple channels for different spatial resolutions corresponding to each atom type. Sunseri et al.[44] facilitates the use of grid-based molecular representations in DL, generating 3D arrays of voxelized molecular data compatible with various DL frameworks.
The research we have reviewed indicates that CNN-based networks excel at encoding pixel-based data, like 2D images and 3D grids, understandable to humans. This ability of CNN to efficiently extract local and global information from such data is essential for analyzing molecular behaviors.
4.1.5 Multi-modality
Multi-modal learning, initially prominent in computer vision, is now widely applied in various fields for its ability to handle and integrate different data types. Its key benefit is enhancing model robustness by using complementary data sources. This approach has gained traction in molecular property prediction.
Due to the significant local chemical information contained in fingerprints may assist models to achieve superior results, Cai et al.[98] and Wang et al.[99] termed fingerprints and graph neural networks, which combined and simultaneously learned information from molecular graphs and fingerprints for MPP. Not only fingerprint, Liu et al.[100], MolFM[101], Sun et al.[102], GraSeq[21] and GIT-Mol[103] employ different encoders to process information from SMILES strings and molecular graphs, respectively. Tang et al.[104] encode molecule by using molecular descriptors and fingerprints, molecular graph and SMILES text notation. Liu et al.[70] combine molecular structural data and textual knowledge to enhance molecular comprehension, jointly learning the chemical structures of molecules and textual knowledge. Zhang et al.[105] use molecular mass spectrum as another representation to provide supplement information which is not contained in the graph data. To address neglects 3D stereochemical information, Chen et al.[106] propose an algebraic graph-assisted bidirectional transformer framework by fusing SMILES and algebraic graph representations. By broad learning of many molecular descriptors and fingerprint features, MolMap[107] was developed for mapping these molecular descriptors and fingerprint features into robust two-dimensional feature maps. To integrate the 3D coordinates information, Zhou et al.[66] employ the atom distance matrix as the position encoding. Liu et al.[108] incorporates comprehensive relational data, including distance, angle, and torsion information between atoms, extending beyond the traditional edge-based 1-hop interactions. Wang et al.[109] embed both molecular graphs and sequences, then create a joint embedding space alongside modality-specific spaces to ensure that the multi-modal data maintains both its distinctive characteristics and a consistent representation across different modalities.
In conclusion, the above work underscores the effectiveness of multi-modal learning in the context of MPP. This approach facilitates the seamless integration of various molecular data formats, including sequences, graph data types, and molecular descriptors. By amalgamating these diverse sources of information, multi-modal learning provides a richer and more nuanced understanding of molecular properties, which is essential for achieving accurate predictions.
4.2 Training Strategy
In this section, we introduce all approaches used to train DL models. While supervised learning has been traditionally predominant, its reliance on scarce labeled data presents limitations.
Figure 6: Training strategy: The high cost of experiments often results in a scarcity of labeled data, leading to challenges like overfitting and poor generalization. To address this, our figure illustrates various advanced training strategies. a). Self-supervised learning, such as contrastive learning and masking recovery, utilizes unlabeled data for pretraining encoders, enabling them to learn molecular domain knowledge effectively. b). Semi-supervised learning methods like pseudo-labeling and co-training leverage both labeled and unlabeled data, enhancing the encoder’s understanding of data distribution. c). Transfer learning strategy capitalizes on pre-trained models from source tasks to boost performance on target tasks. d). Multi-task learning approaches combine related datasets to predict multiple properties simultaneously, benefiting from the relational aspects of molecular data.
To circumvent this, recent approaches have shifted towards unsupervised, self-supervised, and semi-supervised learning methods, capitalizing on the abundance of unlabeled data. Transfer learning is also employed to utilize models pre-trained on unrelated data, enhancing the model’s performance on specific tasks. Additionally, multi-task learning strategies are adopted to leverage related labeled data, further refining the model’s accuracy in predicting molecular properties. As Figure 6 shown, the details of training strategy are as followings.
4.2.1 Self-supervised Learning
Self-supervised learning, widely used in NLP[110, 111], utilizes unlabeled data to extract prior knowledge, proving effective in addressing labeled data scarcity. This method empowers models to learn comprehensive representations from abundant unlabeled data, enhancing their learning capabilities and insight extraction.
Inspired by NLP, Wang et al.[87], Chithrananda et al.[67], Zhang et al.[112], Ahmad et al.[113] and Irwin et al.[114] employ masked language modeling(MLM) on large scale unlabeled data to generate context-sensitive representation, treating SMILES as natural language. Ma et al.[3] use auto-encoder strategy in pretrain stage, first convert SMILES to a vector representation and then reconstructed representation back to SMILES to update the network. Furthermore, Guo et al.[21] fusion the molecular graph and SMILES representation to recontruct the SMILES. Except using SMILES as input, Yuksel et al.[68] employs MLM in SELFIES representations in order to obtain their concise, flexible, and meaningful representations. To let nodes appearing in similar structural contexts to nearby embeddings, Hu et al.[115] propose a context prediction task by using subgraphs to predict their surrounding graph structures.
{forest}
for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center,, text width=2.7cm, s sep=6pt, calign=edge midpoint, text=black, , [ Training
Strategy , fill=gray!45, text=black, parent [ Self-Supervised Learning, text=black, rnn [ Contrastive Learning: Multi-View, rnn_more_cl [ GraphMVP [18]; DVMP [116]; Zhu et al.[20], rnn_work_cl ] ] [ Contrastive Learning: Domain knowledge Boost, rnn_more_cl [ KANO [23]; MoCL [117]; iMolCLR [118], rnn_work_cl ] ] [ Contrastive Learning: Masking Strategy, rnn_more_cl [ MolCLR [22]; GraphCL [119]; ImageMol [15], rnn_work_cl ] ] [ Encoder-Recovery/Prediction, rnn_more_ot [ KPGT [24]; Mole-BERT [120]; K-BERT [121], rnn_work_ot ] ] [ Substructure Enhance, rnn_more_ot [ SME [122]; FragCL [123]; HiMol [73], rnn_work_ot ] ] ] [ Semi-Supervised Learning, text=black, gnn [ Consistency Regularization, gnn_more [ InfoGraph* [26]; DropConn [27], gnn_work ] ] [ Pseudo Label, gnn_more [ ASGN [25]; InstructBio [124], gnn_work ] ] ] [ Transfer Learning, text=black, transformer [ Property-Molecule Relation Enhance, transformer_more [ Meta-GAT [125]; GS-Meta [30], transformer_work ] ] ] [ Multi-Task Learning, text=black, cnn [ Biswas et al. [31], cnn_work ] ] ]
Figure 7: The training strategy summary.
To address GNN oversmoothing and encourage latent node diversity, Godwin et al.[126] employ denoise technique in which they corrupt the input graph with noise, and add a noise correcting node-level loss. Zeng et al.[15] implemented an auto-encoder for molecular image reconstruction, using a discriminator to distinguish between real and fake molecular images. To expand atom vocabulary, Xia et al.[120] use a context-aware tokenizer to encode atom attributes into meaningful discrete codes, then randomly masking and recovering these codes to efficiently pretrain their encoder. Intrinsically, for molecules, a more natural representation is based on their 3D geometric structures, which largely determine the corresponding physical and chemical properties. To overcome the challenge of attaining the coordinate denoising objective, Liu et al.[127] employ an SE(3)-invariant score matching strategy to successfully transform such objective into the denoising of pairwise atomic distances. To capture the anisotropic characteristic of molecules, Feng et al.[128] propose a novel hybrid noise strategy, including noises on both dihedral angel and coordinate, and also decouple the two types of noise and design a novel fractional denoising method, which only denoises the latter coordinate part. For effectively learning 3D spatial representation, Zhou et al.[66] employ 3D position recovery and masked atom prediction as pretrain task. Further more, Jiao et al.[129] exploit the Riemann-Gaussian distribution to ensure the loss to be E(3)-invariant, enabling more robustness. To guild by the molecular domain knowledge and extract chemical information like chemists, Li et al.[9, 24] leverages the molecular descriptors and fingerprints, which serves as the semantics lost in the masked graph to guide the prediction of the masked nodes, thus making the model capture the abundant structural and semantic information from large-scale unlabeled molecules. Wu et al.[121] proposed atom property prediction to discern finer differences between atoms, and MACCS fingerprints prediction, enabling their model to extract and learn predefined molecular features. Gao et al.[130] use atom charges and 3D geometries as inputs, with molecular energies as the target labels, aiming to effectively leverage energy information for enhanced molecular analysis. To optimize multi-task integration and avoid ineffective transfer, Wang et al.[wang2023automated] introduce a fusion strategy that utilizes a surrogate metric based on the total energy of all atoms in a molecule during the pre-training stage. Zang et al.[73] designs three generative tasks that predict bond links, atom types, and bond types with the atom representations and designs two predictive tasks that predict the number of atoms and bonds with the molecule representation. Zeng et al.[131] and Broberg et al.[132] have developed methods to predict the product molecular SMILES based on the reactant molecular embedding. This approach allows for the extraction of chemical information from chemical reactions, providing insights into the molecular transformations involved in the reaction process.
Contrastive learning, a method distinguishing positive and negative molecule pairs, has become a key strategy in encoder pretraining for its ability to enhance molecular structure discernment. This technique is extensively utilized in numerous studies, making it a cornerstone for improving molecular structure recognition in various models. For the SMILES augmentation, Wu et al.[52] and Zhang et al.[133] implemented SMILES enumeration, a technique that varies starting atoms and traversal orders to represent a molecule with different SMILES, thereby uncovering more intricate patterns from complex SMILES structures. Wu et al.[121] and Abdel et al.[134] utilized SMILES permutation as a data augmentation technique, involving the rearrangement of atoms in a SMILES string to create different representations without altering the underlying molecular structure. For molecular graph augmentation, techniques like node dropping, edge perturbation, attribute masking, and subgraph masking are commonly used[119, 22, 135, 136]. However, these random masking methods may not effectively guide the encoder to identify the most crucial chemical information, and might result in the creation of less accurate positive and negative molecule pairs for the training process. To capture important molecular structure and higher order semantic information, Liu et al.[137] adopted the graph attention network as the molecular graph encoder, and leveraged the learned attention weights as masking guidance to generate molecular augmentation graphs. Lin et al.[138] first models the underlying semantic structure of the graph data via clustering semantically similar graphs to select the positive and negative pair and then reweights its negative samples based on the distance between their prototypes and the query prototype such that those negatives having moderate prototype distance enjoy relatively large weights. Cui et al.[139] utilize the GNN encoder and its momentum-update version[140] to generate positive samples at the representation level, and select the negative pairs by the semantic importance of nodes, which is calculated by eigenvector centrality iteration[141]. Wang et al.[142] employ a generative probabilistic model to learn molecular graph structures for topology augmentations and simultaneously develop feature selectors to mask less critical atom features, thus generating effective attribute-level augmentations. To gain deeper insights into chemical information, many researchers incorporate domain knowledge into their contrastive learning approaches. By using backbone and side-chain information, Liu et al.[143] employ side-chain repetition, side-chain generation, backbone disruption, and backbone disruption + side-chain deletion strategy to generate hard positive, soft positive, soft negative and hard negative samples, respectively. Sun et al.[117] replaced a valid substructure by a bioisostere that introduces variation without altering the molecular properties too much, and treat them as positive pairs. Also they optimize the similarity of molecule pairs embedding to be close to the similarity of their ECFP. To avoid faulty negative pairs, Wang et al.[118] mitigate negative contrastive instances by considering ECFP similarities between molecule pairs. Wang et al.[144] calculate the weight vector using the self-attention mechanism to determine the selection probability of each character in SMILES and generate positive samples using three masked strategies: roulette masking, top masking, and random masking. To maintain semantics between conformers, Moon et al.[145] randomly selects molecules from the conformer pool instead of selecting the most stable molecules to learn the 3D structure abundantly. Kuang te al.[146] consider conformations with the same SMILES as positive pairs and the opposites as negative pairs, while keeping the weight to indicate the 3D conformation descriptor and fingerprint similarity. Knowledge graph (KG) is a semantic network composed of entities and their relations in the real world.[147] Hua et al.[148] use the atoms in SMILES as indices to query the embedding matrix to obtain entity and relation embeddings. For the entity and relation vectors of different atoms, they obtain the entity and relation embeddings of the SMILES through linear mapping, and finally concatenate the two vectors to obtain the final embedding representation. Fang et al.[149] first construct a Chemical Element KG based on periodic table of elements, to describes the relations between elements and their basic chemical attributes, Furthermore, they[23] construct another chemical Element KG based on the periodic table and Wikipedia pages to summarize the basic knowledge of elements and the closely related substructure. Those KG offers a comprehensive and standardized view from a chemical element perspective, and help to augment the original molecular graph with the guidance of KG.
In the realm of MPP, a deep understanding of molecular substructures is increasingly recognized as crucial. Many recent studies leverage this domain knowledge to effectively identify and analyze important substructural information, significantly enhancing the understanding of molecular behavior. Xu et al.[150] aimed to preserve local similarities between graph instances by aligning embeddings of related subgraphs and differentiating these from unrelated pairs. They also implemented hierarchical prototypes to represent the latent distribution of graph datasets, enhancing data likelihood with respect to both GNN parameters and these hierarchical structures. Xu et al.[151] aim to capture both intra- and inter-space relationships in group representations and also introduced an attention-based mechanism for generating groups, aimed at identifying diverse molecular substructures effectively. Wang et al.[118] employ BRICS to decompose different substructures which are considered as contrastive negative pairs. Shen et al.[152] and Rong et al.[11] use fingerprints as self-generated labels, Given a graph, its network motifs like chemical functional groups or fragments, which are self-labeled by whether or not they exist in the graph. To learn the local semantics, Luo et al.[153] use graph clustering techniques to partition each whole graph into several subgraphs while preserving as much semantic information as possible, and treat the molecular graph and the clustering graph as postive pair. Benjamin et al.[154] extract substructure information by setting the junction tree(through a tree decomposition algorithm) reconstruction and fingerprint prediction task. To analyze molecular GNN strictly in terms of chemically meaningful fragments, Wu et al.[122] identifies the most crucial set of substructures(BRICS, Murckoand functional groups) in a molecule that are responsible for a model’s prediction. HeGCL[155] introduce the meta-path view that provides semantic information, and encodes graph embeddings by maximizing mutual information between global and semantic representations obtained from the outline and meta-path view, respectively. Hierarchical Molecular Graph is a usually way to extract the substructure molecular representations. Zhu et al.[77] extracts hierarchical information by utilizing co-representation learning of molecular graphs and chemically synthesizable BRICS fragments, and also uses a feature-wise attention block to adaptively recalibrate atomic features after the message passing phase. Kim et al.[123] construct a bag of fragments from a molecule through fragmentation, treating a complete or incomplete bag as a positive or negative view of the original molecule, respectively. Xie et al.[156] proposed a fragment-based molecular graph (FMG) to represent the topological relationship between chemistry-aware substructures within a molecule. They then pre-train it on a fragment level using contrastive learning with well-designed hard negative pairs to extract node representations in FMGs. Xie et al.[156] employ contrastive learning to train fragment encoder, and treat a fragment graph and its associated SMILES string as positive pair, also treat an extremely similar fragment graph serves as a hard negative sample generated by a hard negative construction module. And then encoder the fragment-based molecular graph to extract molecular representation. Xie et al.[156] used contrastive learning for fragment encoder training, pairing a fragment graph with its SMILES string as positive, and pairing a similar fragment graph as negative which generated by a negative construction module, and then applied to fragment-based molecular graphs, enhances molecular representation extraction. Ji et al.[157] decompose the molecular graph using a more reasonable method to construct the fragment graph. They select positive/negative pairs based on similarities between two-level molecule pairs and employ a contrastive loss function, as proposed by Hadsell et al., to pretrain the encoder.
Diverse data formats have been shown to be crucial for MPP, and the multi-modal approach, merging these formats, enhances prediction accuracy by offering a holistic view of molecules. This technique, increasingly adopted in research, combines different data types for a more detailed molecular analysis. To leverage two popular molecular representations and augmentations for each modality, Pinheiro et al.[158], Zhang et al.[159], Zhu et al.[116] and Sun et al.[102] exploit two molecular representations that can be easily acquired from chemical space: the SMILES string and the molecular graph, and then make them as positive pairs. Li et al.[19] utilized self-supervised learning by exploiting the relationship and consistency between 2D topological and 3D geometric structures of molecules. Additionally, Liu et al.[18] applied a generative self-supervised learning approach that focuses on intra-data knowledge, reconstructing key features at the individual data point level to enhance the understanding of molecular structures. 3D Infomax[160] maximized the mutual information between learned 3D summary vectors and the representations of a GNN. Zhu et al.[20] implemented a multifaceted pretraining strategy involving the reconstruction of masked atoms and coordinates, generating 3D conformations based on 2D graphs, and creating 2D graphs from 3D conformations. Kim et al.[123] focused on extracting explicit 3D geometric information by proposing a solution for predicting torsional angles between adjacent molecular fragments, thereby enhancing the depth and accuracy of 3D molecular analysis. Zhu et al.[161] aimed to integrate multiple molecular feature views, including 2D and 3D graphs, Morgan fingerprints, and SMILES strings, ensuring cohesive embedding consistency between these diverse representations for a more unified molecular analysis. The reviewed works show that self-supervised learning, particularly through methods like encoder-recovery and contrastive learning, effectively utilizes unlabeled data to improve model generalization in MPP. These methods excel in learning prior knowledge through various pretraining tasks, allowing for integration of multi-modal data and domain knowledge. This approach significantly enhances the adaptability and performance of models in molecular property prediction scenarios.
4.2.2 Semi-supervised Learning
Semi-supervised learning effectively alleviates the scarcity of labeled molecular data in fields like MPP. By blending a small subset of labeled data with a larger pool of unlabeled data, it bridges the gap between fully supervised and unsupervised learning methods. Consistency regularization is based on the idea that applying realistic perturbations to unlabeled data should not significantly alter predictions, ensuring stability and reliability in the learning process. InfoGraph*[26] employ Mean-Teacher method[162] to maximizes the mutual information between unsupervised graph representations and the representations learned by existing supervised methods in semi-supervised scenarios. Chen et al.[163] predict chemical toxicity and trained the network by the Mean Teacher SSL algorithm, which update the weights in teacher model by applying the Exponential Moving Average. Zhang et al.[27] propose a data augmentation which constructing new adjacency matrix and randomly masking the edges, and calculate the average of all data augmentation distributions and then employ MixMatch[164] label guessing and sharpening method to minimize entropy and accurately guess labels based on the label distribution center.
Proxy-label strategy, assigning temporary labels to unlabeled data, expand the training dataset when labeled data is limited. This approach enhances the model’s learning process, with the proxy labels being iteratively refined for improved accuracy and generalization. ASGN[25] combines a teacher-student framework where the teacher model uses a semi-supervised approach to learn a general representation of molecular graphs and employ an active learning strategy selects informative molecules for labeling. Yu et al.[165] have developed a semi-supervised drug embedding model that combines unsupervised learning from the chemical structures of drugs and drug-like molecules with supervised learning based on hierarchical relations from an expert-crafted drug hierarchy. This approach ensures a robust and comprehensive representation of drug properties. Ma et al.[166] employ teacher-student framework, which use several epochs as a iteration, updating teacher model by the best student model. As the cross-entropy (CE) loss function is not proved to be robust to label noise during the training, they employ generalized CE[167] loss to boost the self-training. To address data imbalance, Liu et al.[168] analyze the distribution of imbalanced annotated data and identify label ranges needing adjustment, and then use high-quality pseudo-labels create graph examples to augment under-represented areas, striving for an ideal balance in training data. Wu et al.[124] introduces an instructor model to provide the confidence ratios as the measurement of pseudo-labels’ reliability. These confidence scores then guide the target model to pay distinct attention to different data points, avoiding the over-reliance on labeled data and the negative influence of incorrect pseudo-annotations.
This approach not only enhances model performance by utilizing the comprehensive information available in unlabeled data but also addresses the challenge of acquiring extensive labeled datasets, a common issue in MPP.
4.2.3 Transfer Learning
Transfer learning strategies, widely adopted in various fields to address data scarcity, focus on enhancing prediction performance for tasks with limited data.[169, 170, 171] These strategies involve transferring knowledge from a data-rich source task to improve learning in a data-scarce target task. Recently, there has been a significant increase in methods employing transfer learning, showcasing its growing importance and application across different domains.
Sun et al.[28] enhanced chemical and physiological property predictions by applying transfer learning, integrating insights from physics and physical chemistry to improve training outcomes. Li et al.[172] developed a framework for accurately estimating task similarity, which, as demonstrated in comprehensive tests, provides valuable guidance for enhancing the prediction performance of transfer learning in molecular property analysis.
Meta-learning, focusing on rapid adaptation to new tasks with minimal data, is effective in addressing the lack of labeled molecular data. Many recent works[173, 174, 175, 176, 177] based on Model-Agnostic Meta-Learning (MAML), enabling rapid adaptation and learning in data-limited scenarios. To effectively utilize correlations of molecules and properties, Lv et al.[125] construct a molecule-property relation Graph, where nodes represent molecules and properties connected by property labels, and then redefine a meta-learning episode as a subgraph within it, containing a target property node along with related molecule and auxiliary property nodes. Chen et al.[29] developed ADKF-IFT, a model that separately trains a subset of parameters with meta-learning loss and adapts others using maximum marginal likelihood for each task. This method, unlike previous ones using a single loss for all parameters, effectively utilizes meta-learning’s regularization to prevent overfitting. MTA[178] is mainly conducting task augmentations by generating new labeled samples through retrieving highly relevant motifs from a pre-defined motif vocabulary as an external memory. To utilize many-to-many correlations of molecules and properties, Zhuang et al.[30] construct a Molecule-Property relation Graph(MPG), then reformulate an episode in meta-learning as a subgraph of the MPG, and then schedule the subgraph sampling process with a contrastive loss function, which considers the consistency and discrimination of subgraphs. Guo et al.[179] developed a model where the importance of different property prediction tasks in few-shot learning is gauged using a self-attentive task weight, calculated by averaging molecular embeddings from each task’s query set, to represent task significance. Wang et al.[180] propose a property-aware embedding function for context-based molecular adaptation and an adaptive relation graph module for molecular relation and embedding refinement, and then employ selective meta-learning strategy for task-specific parameter updates, effectively harmonizing shared knowledge and unique aspects in property prediction tasks. Yao et al.[181] picked out some molecules sharing common properties and use multiple property-aware graph neural networks to extract molecular representation, then employ the Spearman’s correlation to built property-aware matrix. In the few-shot MPP task, the meta-learning strategy is adopted to learn common prediction knowledge from the meta-training categories.
In conclusion, transfer learning has gained popularity for its ability to enhance model generalization in scenarios with limited labeled data. This method is particularly effective in exploiting the relationships between molecules and properties, identifying shared information such as the role of molecular substructures across different tasks, which is crucial for developing more informed and accurate predictive models.
4.2.4 Multi-task Learning
Multi-task learning is a machine learning approach where a model is trained on multiple related tasks simultaneously, rather than training on each task independently. This strategy leverages the commonalities and differences across tasks, allowing the model to learn more generalizable features.
Ma et al.[3] establishing a multi-label supervised model on a combined dataset with missing labels. the input to prediction network is a data matrix with multiple property label information, which can be an original dataset collected from specialized experiments. Tan et al.[32] are putting emphasis on the stacked single-target approach to make the multi-target regression and classification by various single-target base models. Biswas et al.[31] employed a multitask training method for a single model to predict critical properties and acentric factors, while also adjusting target weights in the loss function to correct data imbalance.
These works we’ve reviewed show that multi-task learning is highly effective in MPP, as it capitalizes on the interrelation of various molecular properties. This enhances a model’s capacity to simultaneously predict multiple properties, a particularly valuable trait when dealing with the challenge of limited labeled data.
5 Evaluation and Benchmark
There are a variety of benchmarks available for MPP. Notable among these are Therapeutics Data Commons, MoleculeACE[182], ADMETlab[183], USPTO-MIT[184], and MoleculeNet[33]. Each of these benchmarks offers unique datasets and challenges, allowing for a comprehensive evaluation of different models. This section should detail how models perform across these benchmarks, using a range of metrics to assess accuracy, robustness, and the efficacy of incorporating domain knowledge and multimodal data. By comparing model performances on these diverse benchmarks, we can better understand their capabilities and limitations in the complex field of MPP.
a range of benchmarks were utilized for evaluating Molecular Property Prediction (MPP), including the comprehensive MoleculeNet[33], alongside others such as MoleculeACE[182], ADMETlab[183], and USPTO-MIT[184]. Each benchmark, with MoleculeNet as a primary example, offers unique datasets and challenges, facilitating an exhaustive assessment of various models. Our analysis focuses on model performance across these benchmarks, employing metrics that gauge accuracy, robustness, and the effectiveness of integrating domain knowledge and multimodal data. This approach allows for a nuanced understanding of the strengths and weaknesses of these models in the intricate domain of MPP.
MoleculeNet, a frequently used benchmark in MPP, offers a diverse range of datasets categorized into four groups: Quantum Mechanics, Physical Chemistry, Biophysics, and Physiology. Each group provides specialized datasets to assess different aspects of molecular properties: Quantum Mechanics: Datasets in this group are centered around electronic properties derived from quantum mechanical calculations. Physical Chemistry: These datasets focus on physical and chemical properties of molecules, including solubility and lipophilicity. Biophysics: This category includes datasets related to biological interactions and processes, such as protein-ligand binding affinities. Physiology: Datasets here pertain to organism-level effects, like toxicity and drug efficacy. Evaluating models across these diverse datasets from MoleculeNet allows for a comprehensive assessment of their predictive capabilities in various aspects of MPP.
Consistent with prior studies, we adopt the area under the receiver operating characteristic curve (ROC-AUC) as the evaluation metric for classification datasets, which is a widely used metric for assessing the performance of binary classification tasks. For the regression datasets, we utilize root-mean-squared error (RMSE) as the evaluation metric. It’s important to note that many studies in this field adopt either random or scaffold splits for dividing their datasets, though not uniformly. A random split involves randomly dividing the dataset into training, validation, and test sets, regardless of molecular structures. On the other hand, a scaffold split organizes molecules based on their core chemical scaffolds, ensuring that the model is tested on chemically distinct molecules from those it was trained on, providing a more stringent test of its generalization ability. The choice between these splitting methods can significantly affect the outcomes and interpretations of model performance evaluations.
6 Discussion
6.1 Domain Knowledge Integration
Table 1: Comparison of DL methods for MPP classification tasks with substructure domain knowledge in MoleculeNet. This table contrasts various models, focusing on classification (ROC-AUC %) tasks. Each model is evaluated with and without substructure information, as indicated by original and ablation study rows. The ’-’ symbol marks the absence of data for some datasets, while ’avg. imp.’ shows the average performance improvement due to substructure information integration.
Model Splitting Classification (ROC-AUC % higher is better ↑) avg. imp.
BBBP Tox21 ToxCast SIDER ClinTox BACE HIV
# Molecules \ 2039 7831 8597 1427 1478 1513 41127
# Tasks \ 1 12 617 27 2 1 1
MoLGNN [18] (MoLGNN, GINVAE only) random 88.9 -- -- 63.6 94.2 87.4 78.0 1.09%
89.2 -- -- 61.7 93.7 87.1 76.3
HiGNN [77] (HiGNN, w/o HI) random 93.2 85.6 -- 65.1 93.0 89.0 -- 0.25%
93.0 85.2 -- 65.4 92.6 88.7 --
MISU [154] (MISU, w/o JTVAE) scaffold 66.7 76.3 62.8 59.7 78.0 70.5 -- 1.93%
65.9 76.2 62.3 58.4 76.1 67.1 --
CAFE [156] (CAFE-MPP, Only Graphormer) random 96.5 80.5 -- 65.8 98.2 93.9 -- 3.93%
93.6 79.3 -- 61.8 94.3 89.1 --
iMolCLR [118] (iMolCLR, MolCLR) scaffold 76.4 79.9 73.6 69.9 95.4 88.5 80.8 1.39%
73.6 79.8 72.7 68.0 93.2 89.0 80.6
Table 2: Comparison of DL methods for MPP regression tasks with substructure domain knowledge in MoleculeNet. This table contrasts various models, focusing on regression (RMSE) tasks. Each model is evaluated with and without substructure information, as indicated by original and ablation study rows. The ’-’ symbol marks the absence of data for some datasets, while ’avg. imp.’ shows the average performance improvement due to substructure information integration.
Model Splitting Regression (RMSE, lower is better ↓) avg. imp.
ESOL FreeSolv Lipo QM7 QM8
# Molecules \ 1128 642 4200 6830 21786
# Tasks \ 1 1 1 1 12
HiGNN [77] (HiGNN, w/o HI) random 0.532 0.915 0.549 -- -- 2.78%
0.536 0.941 0.575 -- --
CAFE [156] (CAFE-MPP, Only Graphormer) random 0.687 1.276 0.684 43.75 0.0141 1.37%
0.782 1.303 0.718 40.69 0.0138
iMolCLR [118] (iMolCLR, MolCLR) scaffold 1.130 2.090 0.640 66.30 0.0170 7.78%
1.110 2.200 0.650 87.2 0.0174
This part aims to analyze the contribution of domain knowledge for MPP, as the model input. It is divided into 3 part: atom-bond property, molecular structure, and molecular property relation.
As more and more work utilized the atom and bond property, the effiency of MPP is improved, but whether more atom and bond property integrate in model input? Wojtuch et al.[185] analyzed the impact of atomic features in graph convolutional neural networks, comparing twelve hand-crafted and four literature-based feature combinations. Findings indicate that feature importance is task-specific and linked to their prevalence in the dataset. Reducing less frequent or redundant features, such as formal charges or aromaticity, improves performance. These insights also apply to advanced models like graph transformers, though optimal feature selection varies by model.
Increasingly, molecular structure information is being incorporated into MPP, with several studies leveraging it to derive coarse-grained molecular insights from hierarchical graphs. Recent method like MoLGNN[18], HiGNN[77], MISU[154], CAFE[156], and iMolCLR[118] have used molecular substructure knowledge, such as BRICS or functional groups, to construct hierarchical graphs treating fragments as nodes. These methods have shown improved results over those not using substructure information. As Table 1 and Table 2 demonstrate, the ablation studies reveal a notable enhancement in methodology efficacy when fragment or functional group information is integrated. Specifically, there is a 3.98% improvement observed in regression tasks and a 1.72% improvement in classification tasks, confirming the significant impact of incorporating substructure domain knowledge in these deep learning models.
Identifying a fundamental set of properties for molecular prediction tasks is crucial for future research. Many studies, including multi-task learning methods, have shown that fundamental molecular properties can enhance other prediction tasks. For instance, Sun et al.[28] improved the training of chemical and physiological property predictors by incorporating related physics property prediction tasks. Additionally, Biswas et al.[31] demonstrated the significance of critical properties and acentric factors, along with four phase change properties as auxiliary targets.
6.2 Multi-modal Data Utilization
Table 3: Comparison of multi-modal learning methods in MoleculeNet. The ‘Methods’ column specifies the learning strategy: ‘T’ denotes the use of textual data, ‘2d’ and ‘3d’ refer to the use of 2D and 3D molecular graphs, ‘CL’ indicates contrastive learning, ‘Add’ represents the direct addition fusion method, and ‘CA’ stands for cross-attention fusion. The columns—textual data, graph2d, and graph3d—highlight the types of input data utilized by the models, with the possibility of multiple selections. The best results are emboldened, and the second-best results are highlighted in red.
Model Methods T 2d 3d BBBP Tox21 ToxCast SIDER ClinTox BACE Average
KV-PLM [186] \ 72.0 70.0 55.0 59.8 89.2 78.5 70.8
GIN [187] \ 65.4 74.9 61.6 58.0 58.8 72.6 65.2
GraphMVP [18] (GIN, SchNet) CL(2d, 3d), Add 68.5 74.5 62.7 62.3 79.0 76.8 71.7
72.4 74.4 63.1 63.9 77.5 81.2 72.1
MoMu-S [188] CL(S, 2d) 70.5 75.6 63.4 60.5 79.9 76.7 71.1
MoMu-K [188] 70.1 75.6 63.0 60.4 77.4 77.1 70.6
MoleculeSTM [70] (MegaMolBART, GIN) 70.8 75.7 65.2 63.7 86.6 82.0 74.0
70.0 76.9 65.1 61.0 92.5 80.8 74.4
MolFM [101] (KV-PLM, GIN) CL(S, 2d), CA 72.2 76.6 64.2 63.2 78.6 82.6 72.9
72.9 77.2 64.4 64.2 79.7 83.9 73.7
GIT-Mol [103] (SciBERT, MoMu-S) 71.9 73.9 62.1 60.1 83.5 68.4 70.0
71.1 75.4 65.3 58.2 78.9 65.8 69.1
73.9 75.9 66.8 63.4 88.3 81.1 74.9
MolLM [189] CL(S, 2d), CL(S, 3d) 75.7 80.0 68.2 71.0 91.1 84.1 78.4
This part aims to analyze the contribution of different modalities in multi-modal models for MPP. It delves into understanding how individual and combined modalities affect prediction performance. Data from various studies are collated and analyzed, emphasizing the contribution of different modalities to MPP tasks, with models in Table 3. Apart from ClinTox, there is uniformity in the predictive prowess displayed by all models across a spectrum of tasks. Nonetheless, ClinTox predictions are prone to biases in transformer-based models due to data distribution peculiarities, which result in polarized predictions. Graph-based models demonstrate a reduction in such bias, albeit with a compromised performance in ClinTox. From an input modality perspective, 3D graphical modalities exhibit superior performance. Excluding ClinTox, the performance of SMILES and graph2d is on par, with graph2d holding a slender advantage. In examining pre-training strategies, contrastive learning clearly demonstrates significant benefits. However, the integration of cross-attention might inadvertently reduce the impact of the singular modalities. Nonetheless, the strategic implementation of cross-attention promotes an effective fusion of SMILES and graph2d, resulting in combined vectors that outperform the individual modalities. In cases where modality fusion is employed, an analysis focused on the contributory significance of each data modality to the tasks reveals an interesting pattern: The standalone modalities of SMILES and graph2d yield an average performance of around 71.6%. In comparison, their fusion (SMILES+graph2d) escalates the average performance to about 74.6%. This increment attests to the potency of modality fusion, enhancing performance by an estimated 4.2%.
In conclusion, the 3D graphical modality, enriched with voluminous molecular data, is adjudged the top performer, while SMILES and graph2d exhibit comparable efficacies. In the clinTox task, addressing model bias presents a significant research opportunity. Modality fusion methods demonstrate advantages, suggesting substantial exploration potential in the ways modalities are fused.
7 Conclusion
In this paper, we discuss the significant role of multi-modal data and domain knowledge in enhancing molecular property prediction through DL methods. We explored various molecular data formats and domain knowledge, crucial in understanding molecular complexities. Our review of recent encoder architectures and training strategies highlighted how integrating domain knowledge and multi-modal data advances these models. By benchmarking prominent works, we provided a comparative analysis of their effectiveness. Ultimately, our discussion revealed the profound impact of domain knowledge and multi-modal data in DL approaches, marking a transformative advancement in drug discovery and computational molecular analysis.